# Why Rectified Power Unit Networks Fail and How to Improve It: An Effective Field Theory Perspective

### v4

@@ 5.2 Experiments on Synthetic Data
Case of Physics-Informed Neural Networks (PINNs) Next, we compare and contrast neural networks constructed with different activation functions using the physics-informed neural networks (PINNs) (Raissi19) technique, which solves PDE problems by incorporating the PDE itself into the loss function. The PDE used in this experiment is the Burgers equation, given as:
@@
The results of the experiment are presented in Fig. 15, 16, and 17. Interestingly, unlike the previous cases with polynomials and differentiable functions, the GELU activation exhibited the best performance, followed by the MRePU network. The ReLU network, however, failed to achieve meaningful learning. Additional experiments with the Tanh activation showed a performance similar to that of the GELU network. From this experiment, it can be concluded that the MRePU network is capable of achieving meaningful learning even under the highly complex training dynamics of the PINN loss (Wang22, Aditi21) and with a deep architecture consisting of 7 hidden layers (9 layers in total).
@@ 5.3 Experiments on Real-World Data
MNIST Classification First, we conducted experiments on the MNIST dataset, which is fundamental and widely used for verification. The $28\times 28$ input images were flattened into 784-dimensional vectors, and a 4-layer FCN architecture with dimensions $[784,256,256,$ $256,10]$ was employed. In all experiments, training was performed using the Adam optimizer with a learning rate of $10^{-3}$. For the first experiment verifying the criticality condition, a batch size of 512 was used, while in the second experiment comparing the performance of various activation functions, a batch size of 64 was utilized. In the first experiment, to empirically demonstrate that the condition in Eq. (11) serves as a boundary condition for training stability, the dataset was centered and scaled by dividing by $N_{K}$ (ranging from 0.1 to 2.0 with an interval of 0.1). Similarly, the hyperparameter $C_{W}$ was varied from 0.1 to 2.0 with an interval of 0.1. We performed training for 5 epochs for each case, repeated the process 10 times, and visualized the average test accuracy as a heatmap (Fig. 18).By regarding $C_{W}$ in Eq. (11) as a function $\tilde{C}_{W}(K)$ of $K$ for the case where $p=2$, the criticality condition or phase boundary obtained via fitting, $2.2\tilde{C}_{W}(0.02/N_{K})$, is depicted as a dotted red line in Fig. 18. As shown in the figure, the success of neural network training is highly sensitive to $C_{W}$, and the fitted threshold condition aligns well with the empirical observations. Notably, as $N_{K}$ increases and the data approaches the fixed point $K^{\star}=0$, the sensitivity to $C_{W}$ diminishes.
@@
CIFAR-10 Classification For the CIFAR-10 dataset, we employed the ResNet architecture. For ReLU and GELU, we adhered to the architecture described in (He15b), utilizing the basic building block. The network is constructed by arranging blocks with widths of 64, 128, 256, and 512 in a sequence of 2, 2, 1, and 1 blocks, respectively. Each block consists of two $3\times 3$ convolutional layers with a residual connection added to the input and output. Including the input and output layers, the total depth of the architecture is $1+2(2+2+1+1)=14$ layers.
For MRePU, we modified the residual structure as follows, which is an essential adjustment to maintain the kernel at a constant scale (Roberts22):
@@
where $z^{(l+1)}$ is final output of whole block and $z^{(l)}$ is input of the block $0<\gamma<1$ is a weight for the residual term. Since MRePU, ReLU, and GELU all share $K^{\star}=0$ as a fixed point, it is fair to ensure the kernel scale remains close to zero. Consequently, we normalized the dataset to the following scale:
@@
We utilized a batch size of 64 and the Adam optimizer with a learning rate of $2.5\times 10^{-4}$. The training was conducted for 5 epochs. We performed a total of 25 independent trials for each activation function, and the average test loss and accuracy are reported in Table 6. The moving average of the log-scaled training loss, calculated over a window of 50 iterations, is depicted in the right panel of Fig. 19. While all three activation functions exhibit similar trends, it is observed that, unlike in the MNIST case, the reduction in training loss for MRePU lags slightly behind the others.
@@
Table 6: Mean final training losses and accuracies over 25 trials for real-world tasks for each of the MRePU (order $p=2$), ReLU, and GELU networks. The values in bold represent the lowest test loss values(highest for accuracies) in each column.

### v5

@@ 5.2 Experiments on Synthetic Data
Case of Physics-Informed Neural Networks (PINNs) Next, we compare and contrast neural networks constructed with different activation functions using the physics-informed neural networks (PINNs) (Raissi et al. (2019)) technique, which solves PDE problems by incorporating the PDE itself into the loss function. The PDE used in this experiment is the Burgers equation, given as:
@@
The results of the experiment are presented in Fig. 15, 16, and 17. Interestingly, unlike the previous cases with polynomials and differentiable functions, the GELU activation exhibited the best performance, followed by the MRePU network. The ReLU network, however, failed to achieve meaningful learning. Additional experiments with the Tanh activation showed a performance similar to that of the GELU network. From this experiment, it can be concluded that the MRePU network is capable of achieving meaningful learning even under the highly complex training dynamics of the PINN loss (Wang et al. (2022), Krishnapriyan et al. (2021)) and with a deep architecture consisting of 7 hidden layers (9 layers in total).
@@ 5.3 Experiments on Real-World Data
MNIST Classification First, we conducted experiments on the MNIST dataset, which is fundamental and widely used for verification. The $28\times 28$ input images were flattened into 784-dimensional vectors, and a 4-layer FCN architecture with dimensions $[784,256,256,$ $256,10]$ was employed. In all experiments, training was performed using the Adam optimizer with a learning rate of $10^{-3}$. For the first experiment verifying the criticality condition, a batch size of 512 was used, while in the second experiment comparing the performance of various activation functions, a batch size of 64 was utilized. In the first experiment, to empirically demonstrate that the condition in Eq. (11) serves as a boundary condition for training stability, the dataset was centered and scaled by dividing by $N_{K}$ (ranging from 0.1 to 2.0 with an interval of 0.1). Similarly, the hyperparameter $C_{W}$ was varied from 0.1 to 2.0 with an interval of 0.1. We performed training for 5 epochs for each case, repeated the process 10 times, and visualized the average test accuracy as a heatmap (Fig. 18). By regarding $C_{W}$ in Eq. (11) as a function $\tilde{C}_{W}(K)$ of $K$ for the case where $p=2$, the criticality condition or phase boundary obtained via fitting form of function $a\tilde{C}_{W}(b/N_{K})$ into $2.2\tilde{C}_{W}(0.02/N_{K})$, is depicted as a dotted red line in Fig. 18. As shown in the figure, the success of neural network training is highly sensitive to $C_{W}$, and the fitted threshold condition aligns well with the empirical observations. Notably, as $N_{K}$ increases and the data approaches the fixed point $K^{\star}=0$, the sensitivity to $C_{W}$ diminishes.
@@
CIFAR-10 Classification For the CIFAR-10 dataset, we employed the ResNet architecture. For ReLU and GELU, we adhered to the architecture described in (He et al. (2016)), utilizing the basic building block. The network is constructed by arranging blocks with widths of 64, 128, 256, and 512 in a sequence of 2, 2, 1, and 1 blocks, respectively. Each block consists of two $3\times 3$ convolutional layers with a residual connection added to the input and output. Including the input and output layers, the total depth of the architecture is $1+2(2+2+1+1)=14$ layers.
For MRePU, we modified the residual structure as follows, which is an essential adjustment to maintain the kernel at a constant scale (Roberts et al. (2022)):
@@
where $z^{(l+1)}$ is final output of whole block and $z^{(l)}$ is input of the block $0<\gamma<1$ is a weight for the residual term. Since MRePU, ReLU, and GELU all share $K^{\star}=0$ as a fixed point, it is fair to ensure the kernel scale remains close to zero. Consequently, we normalized the dataset to the following scale, where scaling factor in denominator is empirically tuned:
@@
We utilized a batch size of 64 and the Adam optimizer with a learning rate of $2.5\times 10^{-4}$. The training was conducted for 5 epochs. We performed a total of 25 independent trials for each activation function, and the average test loss and accuracy are reported in Table 6. The moving average of the log-scaled training loss, calculated over a window of 50 iterations, is depicted in the right panel of Fig. 19. While all three activation functions exhibit similar trends, it is observed that, unlike in the MNIST case, the reduction in training loss for MRePU lags slightly behind the others. Although the training-loss curves exhibit qualitatively similar trends across activations, MRePU (with $p=2$) yields a noticeably higher test loss and lower test accuracy on CIFAR-10 (Table 6). Nevertheless, the main significance of MRePU lies in enabling stable optimization in deeper and more complex architectures such as ResNets under the proposed criticality condition. In our experiments, such stable training was not observed when using RePU, or when the criticality condition was violated, where optimization frequently became unstable and failed to converge.
@@
Table 6: Mean final test losses and accuracies over 25 trials for real-world tasks for each of the MRePU (order $p=2$), ReLU, and GELU networks. The values in bold represent the lowest test loss values(highest for accuracies) in each column.